OpenStudy (anonymous):

Medal for whoever helps, need quick explanantion, not a whole essay explaining this! m = -6 with x-intercept 12 Point Slope Form:_____________________ Standard Form:_____________________ Slope-Intercept form:_____________________

OpenStudy (anonymous):

Point slope form: y – y1 = m(x – x1) slope intercept form: y=mx+b Standard Form: ax+by=c

OpenStudy (anonymous):

Point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

y basically means y2 and x mean x2

OpenStudy (anonymous):

Hope thats puts you on the right track.